# New Foundations is consistent

### v22

@@ 2.4.2 Tangled webs of cardinals: a suggestion of another approach not followed here
If there is a model of Mac Lane set theory in which there is a tangled web of cardinals $\tau$, then NF is consistent.
@@
Let $\Sigma$ be a finite set of sentences of the language of TST. Let $n$ be larger than any type mentioned in any formula in $\Sigma$. Partition $[\alpha]^{n}$ into compartments in such a way that the compartment that a set $A$ is put into depends on the truth values of the sentences in $\Sigma$ in natural models of TSTn with base type of size $\tau(B)$ where $B\setminus B_{n}=A$. This partition of $[\alpha]^{n}$ into no more than $2^{|\Sigma|}$ compartments has a homogeneous set $H$ of size $n+1$. The natural models of TSTn with base types of size $\tau(H)$ and base types of size $\tau(H_{1})$ have the same truth values for sentences in $\Sigma$, so the model of TST with base type $\tau(H)$ satisfies the restriction of the Ambiguity Scheme to $\Sigma$, so the full Ambiguity Scheme is consistent by compactness, so TST + Ambiguity is consistent so NF is consistent. ∎
Our initial approach to proving our theorem was to attempt a Frankel-Mostowski construction of a model of Mac Lane set theory with a tangled web of cardinals. We do know how to do this, but we believe from recent experience that constructing a model of tangled type theory directly is easier, though tangled type theory is a nastier theory to describe.
We think there is merit in giving a brief description of a situation in a more familiar set theory equivalent to (a strengthening of) the very strange situation in a model of tangled type theory. This section is also useful here because it supports the discussion in the conclusion of one of the unsolved problems which is settled by this paper.
@@ 2.5 An axiomatization of TST with finitely many templates
We discuss a finite axiomization of NF derived from that of Hailperin (it is taken from an implementation of Hailperin’s axiom set of [4] in Metamath ([2]), and there are minor changes from the original formulation), making the important observation that it actually provides us with an axiomatization of TST with finitely many axiom templates (in the sense that each axiom is a type shifted version of one of a finite set of axioms). Notation introduced in this section is not used in the rest of the paper, and nothing in subsequent sections except a brief remark in the last paragraphs of section 4 depends on anything here. This finite axiomatization is however used in the Lean formalization.
@@
$(\forall zw:(z,w)\in R^{-1}\leftrightarrow(w,z)\in R)\wedge(\forall u:u\in R^{-1}\leftrightarrow(\exists zw:(z,w)=u)).$
@@
We define $\iota^{2}``x$ as $(x\times x)\cap V^{\iota}$. This is the image of $x$ under the double singleton operation.
@@
We define $I_{2}(R)$ as the intersection of a set provided by the previous axiom with $V\times V\times V$.
@@
We define $I_{3}(R)$ as the intersection of a set provided by the previous axiom with $V\times V\times V$.
@@
It seems natural to define $I_{1}(R)$ as $\iota^{2}``V\times R$, but this requires no new axiom.
@@
Define $I_{1,n}(R)$ as $\iota^{2(n-1)}``V\times R$: this is correct for prepending all possible initial projections to an $n$-tuple. Then define $I^{1}_{1,n}(R)$ as $I_{1,n}(R)$ and define $I^{m+1}_{1,n}(R)$ as $I_{1,n+m}(I^{m}_{1,n}(R))$: this takes into account the fact that the tuples get longer.
Define $I_{2,n}(R)$ and $I^{1}_{2,n}(R)$ as $I_{2}(R)\cap(V\times\iota^{2(n-3)}(V)\times V)$, and define $I^{m+1}_{2,n}(R)$ as $I_{1,n+m}(I^{m}_{2,n}(R))$: this takes into account the fact that the tuples get longer.
@@
The set $\{(x_{1},\dots,x_{n}):x_{1}=x_{i}\}$ ($i<n$) is obtained as $I_{2,3}^{i-3}(I_{3}(\iota^{2}``V))\cap\tau^{*}$.
@@
The set $\{(x_{1},\ldots,x_{n}):(\exists x_{1}:(x_{1},\ldots,x_{n})\in R)\}$ is representable as $I_{1,n-1}({\tt dom}(R^{-1}))$.
This only allows us to quantify over the lowest numbered variable. This is actually sufficient. We have renamed bound variables so that all bound variables with different binders are distinct and in any subformula $(\exists y:\psi)$ $y$ will have lower index than any variable free in $\psi$. Then the quantified variable might not be $x_{1}$, but if it is $x_{k}$, we can use the representation $I^{k}_{1,n-k}({\tt dom}({\tt rng}^{k-1}(R))^{-1})$: strip off the first $k-1$ variables, which are dummies already quantified over, quantify over $x_{k}$ and put $k$ dummies back, as it were.
@@
where a final bit of computation must be exhibited: $\{(x_{1},\ldots,x_{n}):x_{i}=t\}$ is realized as $I^{i-1}_{1,2}(\{\iota^{2(n-i-1)}(t_{i})\}\times V)\cap\tau^{*}.$
We state for the record that we think this is a bad finite axiomatization of NF, or finite template axiomatization of TST: we think the treatment of $n$-tuples is terribly difficult to work with. But it does work, and the second author chose to verify it, so it merits discussion here. It should be noted that any finite template axiomatization of TST could be used; there is no advantage to this one for the formalization, and there is an oddity, because the one impredicative axiom (the axiom of type lowering) does a lot of extra work, since it is also essential for constructing domains and ranges of functions. In the formalization, the proof of type lowering was in effect divided into the proof of the existence of $\iota^{3}``{\tt TL}(x)$, which is predicative, and the proof of the existence of set unions of sets of singletons.
@@ 3.1 The abstract supertype framework
Foundation in the metatheory ensures a clean construction here. An element $x$ of supertype $t>_{\tau}m$ is always nonempty with $t^{+}$ as an element. The set $t^{+}$ has supertype $u$ as an element for each $u<_{\tau}t$, so $t^{+}$ and so $x$ cannot belong to any supertype $u$ with $u<_{\tau}t$, by foundation: each other element of $x$ belongs to such a supertype. We have shown that all the types are disjoint. The labelling element $t^{+}$ cannot belong to supertype $t$ by foundation, because an element of supertype $t$ must be nonempty and have $t^{+}$ as an element. Further, $t^{+}$ cannot belong to any supertype $v$ with $t<_{\tau}v$, because any element of $v$ contains $v^{+}$ as an element which contains supertype $t$ as an element and any element of supertype $t$ contains $t^{+}$ as an element, so $t^{+}\in v$ would violate foundation in the metatheory.
@@
Let $M$ be a model of TTT (more generally, any structure for the language of TTT in which each object is determined given all of its extensions). Let $\leq_{M}$ be the well-ordering on the types of $M$ and let $m$ be the minimal type of $M$. We will assume as above that $\leq_{\tau}$ is a well-ordering of type labels $t$ with corresponding actual types $\tau_{t}$ of $M$: of course, we could use the actual types of $M$ as type indices, but we preserve generality this way. We also assume that the sets implementing the types of $M$ are disjoint (it is straightforward to transform a model in which the sets implementing the types are not disjoint to one in which they are, without disturbing its theory, by replacing each $x\in\tau_{t}$ with $(x,t)$).
@@
If we have defined $I$ on each type $u<_{\tau}t$, we define $I(x)$, for $x\in\tau_{t}$, as $\bigcup_{u<_{\tau}t}\{I(y):y\in^{M}_{u,t}x\}\cup\{\{\tau^{*}_{u}:u<_{\tau}t\}\}$.
@@
Notice that if $\alpha>_{\tau}\beta$ are supertypes, and $x\in\tau^{*}_{\alpha}$, $x\cap\tau^{*}_{\beta}$ is the extension of $x$ over supertype $\beta$. This will be inherited by a scheme of types $\tau_{\gamma}$ with each $\tau_{\gamma}\subseteq\tau^{*}_{\gamma}$ if an additional condition holds: for $\alpha>_{\tau}\beta$, we will have for $x\in\tau_{\alpha}$ that $x\cap\tau^{*}_{\beta}$ is the extension of $x$ over supertype $\beta$ as already noted: for it to be the extension over type $\beta$ we need the general condition $x\cap\tau^{*}_{\gamma}\subseteq\tau_{\gamma}$ for all $\delta>\gamma$ and $x\in\tau_{\delta}$.
@@ 3.2 Preliminaries of the construction
On each litter $L=L_{\nu,\beta,\gamma}$ define a well-ordering $\leq_{L}$: $(\nu,\beta,\gamma,\alpha)\leq_{L}(\nu,\beta,\gamma,\alpha^{\prime})$ iff $\alpha\leq\alpha^{\prime}$. The strict well-ordering $<_{L}$ is defined in the obvious way. This well-ordering is used in only one place in the paper (theorem 4.18), and its use could easily be avoided, but we find its concreteness appealing.
@@
A near-litter is determined as the symmetric difference of a litter $L$ (there are $\mu$ litters) and a small subset of $\tau_{-1}$. So it is sufficient to show that a set of size $\mu$ has no more than $\mu$ small subsets. As $\mu$ has cofinality at least $\kappa$, each small subset of $\mu$ is bounded, and so is an element of $\bigcup_{\nu<\mu}\mathcal{P}(\nu)$. But as $\mu$ is a strong limit cardinal, each $\mathcal{P}(\nu)$ has size less than $\mu$, so there are only $\mu$ bounded subsets of $\mu$; in particular, there are only $\mu$ small subsets of $\mu$. ∎

### v23

@@ 2.3.1 There are $\alpha$-models for every $\alpha$
For any infinite ordinal $\alpha$ of the metatheory, if there is a model of TTT(U) in which the ordinal $\alpha$ has an implementation $\alpha^{s}$ for each strictly increasing sequence $s$ of types, by which we mean an ordinal in type $s_{4}$ whose elements of type $s_{3}$ correspond to well-orderings of order type $\alpha$ of subsets of type $s_{0}$, using types $s_{1}$ and $s_{2}$ for the intermediate pairing machinery, and which is such that the ordinal $\lambda$ indexing the types is a strong limit cardinal of cofinality $>|\alpha|^{|\alpha|}$, then there is a model of NF(U) in which the ordinal $\alpha$ is implemented, which further satisfies [after neglect of types] any sentence in the language of TST(U) whose translations via all increasing sequences of types are satisfied in the model of TTT(U).
@@
The proof is a modified version of the proof given above for consistency of NF(U), using a more powerful partition principle. We do not claim originality, referring the reader to [10], but our presentation is not the same, and it is adapted to talk of tangled type theory.
The Erdős–Rado theorem tells us that for a natural number $n\geq 1$ and an infinite cardinal $\mu$, any partition of $\exp^{n}(\mu)$ into $\mu$ compartments has a homogeneous set of size $\mu^{+}$. This is not quite the usual formulation, but easily follows from it.
We extend the language of TTT(U) to include constant terms $\alpha^{s}$ and $\beta^{s}$ for each ordinal $\beta<\alpha$ of the metatheory for each sequence $s$ as described above. We similarly extend the language of TST(U) (in this case only needing constants $\alpha^{i}$ and $\beta^{i}$ for $\beta<\alpha$ for each $i\geq 4$), and note that the translation of a sentence of TST(U) with language augmented with these ordinal constants into terms of a strictly increasing sequence of types in TTT(U) is straightforward, and we allow use of these constants in the comprehension axiom of TTT(U) as usual. We further include the Hilbert symbol in our language, with the same restriction stated above that the Hilbert symbol may not appear in instances of comprehension, noting our intention to construct a term model.
We consider partitions of $[\lambda]^{n}$ determined by the set $\Sigma_{n}$ of all closed Hilbert symbols (“closed” meaning the symbol contains no parameters) of the form $(\epsilon\beta^{i}:\beta^{i}<\alpha^{i}\wedge\phi)$ in the language of TST(U) augmented as indicated above and using only types less than $n$, with the compartment in the partition containing $A\in[\lambda]^{n}$ determined by the ordinal values $<\alpha$ [or default value, where we assume the default value is not an ordinal] assigned to all of these terms in our model of TTT(U) when types $0-(n-1)$ are replaced by the types in $A$ in increasing order in each of the Hilbert symbols in $\Sigma_{n}$. Here we are speaking loosely in that we say an ordinal $\gamma<\alpha$ is assigned as value to $(\epsilon\beta^{i}:\beta^{i}<\alpha^{i}\wedge\phi)$ when its value is in fact $\gamma^{i}$: it should be clear that this is unproblematic and easier to say.
There are no more than $|\alpha|$ such Hilbert symbols, and no more than $|\alpha|^{|\alpha|}$ compartments in each of these partitions.
Note that assignment of values to all such Hilbert symbols induces a complete assignment of truth values to sentences of the extended language:
$(\epsilon\beta^{i}:\beta^{i}<\alpha\wedge\beta^{i}<2\wedge(\beta=1\leftrightarrow\phi))$
codes a sentence $\phi$ into $\Sigma_{n}$ if $\phi$ mentions only types below $n>i$, because the assignment of ordinal value to this term corresponds to the assignment of truth value to $\phi$.
The Erdős–Rado theorem allows us to conclude that for any infinite $\mu<\lambda$, we have a homogeneous set of size $\mu^{+}$ for the given partition in $\exp^{n}(\mu)<\lambda$. Further, we can find a single assignment of values to Hilbert symbols as described above which is associated with arbitrarily large homogeneous sets below $\lambda$, because the cofinality of $\lambda$ is greater than $|\alpha|^{|\alpha|}$, the cardinality of the set of compartments in the partition, corresponding precisely to the cardinality of the set of such assignments of values. If we have chosen such an assignment of values for a specific value of $n$, we can further choose such an assignment of values for $n+1$ which extends it: for any $\mu$, choose a homogeneous set for the partition of $[\lambda]^{n}$ realizing the fixed assignment of values and of cardinality $\nu>\exp^{n+1}(\mu)$, then find a homogeneous set of size $\mu^{+}$ for the partition of $[\lambda]^{n+1}$ inside this homogeneous set, which will determine an assignment of values to Hilbert symbols using $n+1$ types extending the fixed one chosen for Hilbert symbols using $n$ types. Thus there are arbitrarily large homogeneous sets for assignments of values to Hilbert symbols using $n+1$ types extending the fixed one for Hilbert symbols using $n$ types. Thus there is a fixed assignment of values to Hilbert symbols using $n+1$ types which extends the given fixed assignment for $n$ types and which is associated with arbitrarily large homogeneous sets, by the already stated considerations of cofinality. Thus we can proceed through $\omega$ steps and find an assignment of values to all Hilbert symbols with the property that its restriction to any fixed number $n$ of types is realized in arbitrarily large homogenous sets below $\lambda$.
Thus we have constructed a complete assignment of values to such Hilbert symbols in the extended TST(U) language, and thus a complete assignment of truth values to all sentences, which is ambiguous [invariant under displacement of types] and consistent with any statement holding at all sequences of types in the model of TTT(U). Thus for example we will get extensionality if the model of TTTU is actually a model of TTT. In fact, for any fixed $n$ all these assignments for Hilbert symbols mentioning only types less than $n$ will be realized at some sequence of types in the model of TTT(U) [verifying that this assignment of values is consistent and in effect a complete description of a model of NF(U)].
Further, this model of NFU contains no ordinal less than what it calls $\alpha$ other than the ordinals which it calls $\beta$ for $\beta<\alpha$ in the metatheory, and the natural well-ordering on these ordinals is realized in the model of NF(U) and corresponds to a well-ordering of order type $\alpha$ in the metatheory. The reason for this is that the model of NFU is a term model, and any term denoting an ordinal less than $\alpha$ can be slightly modified to the exact form used in defining the partitions, and the process will assign that term a concrete value less than $\alpha$, an actual ordinal $\beta<\alpha$ of the metatheory, so the order on the ordinals $<\alpha$ in the model of NF(U) has order type $\alpha$ in the metatheory.
@@ 2.4 An axiomatization of TST with finitely many templates
We discuss a finite axiomatization of NF derived from that of Hailperin (the statement of the axioms is taken from an implementation of Hailperin’s axiom set of [4] in Metamath ([2]), and there are minor changes from the original formulation; the proofs are our own), making the important observation that it actually provides us with an axiomatization of TST with finitely many axiom templates (in the sense that each axiom is a type shifted version of one of a finite set of axioms). Notation introduced in this section is not used in the rest of the paper, and nothing in subsequent sections except a brief remark in the last paragraphs of section 4 depends on anything here. This finite axiomatization is however used in the Lean formalization, which is why we review it in detail here.
@@
$(\forall zw:(z,w)\in R^{-1}\leftrightarrow(w,z)\in R)\wedge(\forall u:u\in R^{-1}\rightarrow(\exists zw:(z,w)=u)).$
@@
definitions:
We define $\iota^{2}``x$ as $(x\times x)\cap 1$. This is the image of $x$ under the double singleton operation.
We define $V^{1}$ as $V$ and $V^{n+1}$ as $\iota^{2(n-1)}``V\times V^{n}$, for each concrete $n$. $V^{n}$ is the set of all $n$-tuples.
@@
We define $I_{2}(x)$ as the intersection of a set provided by the previous axiom with $V\times V\times V$.
@@
We define $I_{3}(x)$ as the intersection of a set provided by the previous axiom with $V\times V\times V$.
@@
It seems natural to define $I_{1}(R)$ as $\iota^{2}``V\times R$; this requires no new axiom.
@@
Define $I_{1,n}(R)$ as $\iota^{2(n-1)}``V\times R$: this is correct for prepending all possible initial projections to an $n$-tuple. Then define $I^{1}_{1,n}(R)$ as $I_{1,n}(R)$ and define $I^{m+1}_{1,n}(R)$ as $I_{1,n+m}(I^{m}_{1,n}(R))$: this takes into account the fact that the tuples get longer. The intention here is that $I^{m}_{1,n}$ represents the result of applying the insertion one operation to an $n$-ary relation $m$ times, yielding an $(n+m)$-ary relation.
Define $I_{2,n}(R)$ and $I^{1}_{2,n}(R)$ as $I_{2}(R)\cap V^{n+1}$ [one effect of the intersection with $V^{n+1}$ is to restrict second projections of tuples in $I_{2}(R)$ to appropriately iterated singletons], and define $I^{m+1}_{2,n}(R)$ as $I_{2,n+m}(I^{m}_{2,n}(R))$: this takes into account the fact that the tuples get longer. The intention here is that $I^{m}_{2,n}$ represents the result of applying the insertion two operation to an $n$-ary relation $m$ times, yielding an $(n+m)$-ary relation.
@@
The set $\{(x_{1},\dots,x_{n}):x_{1}=x_{i}\}$ ($i<n$) is obtained as $I_{2,3}^{i-2}(I_{3}(\iota^{2}``V))\cap\tau^{*}$.
@@
The set $\{(x_{1},\ldots,x_{n}):(\exists x_{1}:(x_{1},\ldots,x_{n})\in R)\}$ is representable as $I_{1,n-1}({\tt rng}(R))$.
This only allows us to quantify over the lowest numbered variable. This is actually sufficient. We have renamed bound variables so that all bound variables with different binders are distinct and in any subformula $(\exists y:\psi)$ $y$ will have lower index than any variable free in $\psi$. Then the quantified variable might not be $x_{1}$, but if it is $x_{k}$, we can use the representation $I^{k}_{1,n-k}({\tt rng}^{k}(R))$: strip off the first $k-1$ variables, which are dummies already quantified over, quantify over $x_{k}$ (stripping off one more variable) and put $k$ dummies back, as it were.
@@
where a final bit of computation must be exhibited: $\{(x_{1},\ldots,x_{n}):x_{i}=t\}$ is realized as $I^{i-1}_{1,2}(\{\iota^{2(n-i-1)}(t)\}\times V)\cap\tau^{*}.$
We state for the record that we think this is a bad finite axiomatization of NF, or finite template axiomatization of TST: we think the treatment of $n$-tuples is terribly difficult to work with. But it does work, and the second author chose to verify the consistency of TTT by verifying that the finite (template) axiomatization is satisfied in the structure for the language of TTT that we define, so it merits discussion here. It should be noted that any finite template axiomatization of TST could be used; there is no advantage to this one for the formalization, and there is an oddity, because the one impredicative axiom (the axiom of type lowering) does a lot of extra work, since it is also essential for constructing domains and ranges of functions. In the formalization, the proof of type lowering was in effect divided into the proof of the existence of $\iota^{3}``{\tt TL}(x)$, which is predicative, and the proof of the existence of set unions of sets of singletons.
@@ 3.1 The abstract supertype framework
Foundation in the metatheory ensures a clean construction here. An element $x$ of supertype $t>_{\tau}m$ is always nonempty with $t^{+}$ as an element. The set $t^{+}$ has supertype $u$ as an element for each $u<_{\tau}t$, so $t^{+}$ and so $x$ cannot belong to any supertype $u$ with $u<_{\tau}t$, by foundation. The labelling element $t^{+}$ cannot belong to supertype $t$ by foundation, because an element of supertype $t$ must be nonempty and have $t^{+}$ as an element. Further, $t^{+}$ cannot belong to any supertype $v$ with $t<_{\tau}v$, because any element of $v$ contains $v^{+}$ as an element which contains supertype $t$ as an element and any element of supertype $t$ contains $t^{+}$ as an element, so $t^{+}\in v$ would violate foundation in the metatheory.
@@
Let $M$ be a model of TTT (more generally, any structure for the language of TTT in which each object outside the base type is determined given all of its extensions). Let $\leq_{M}$ be the well-ordering on the types of $M$ and let $m$ be the minimal type of $M$. We will assume as above that $\leq_{\tau}$ is a well-ordering of type labels $t$ with corresponding actual types $\tau_{t}$ of $M$: of course, we could use the actual types of $M$ as type indices, but we preserve generality this way. We also assume that the sets implementing the types of $M$ are disjoint (it is straightforward to transform a model in which the sets implementing the types are not disjoint to one in which they are, without disturbing its theory, by replacing each $x\in\tau_{t}$ with $(x,t)$).
@@
If we have defined $I$ on each type $u<_{\tau}t$, we define $I(x)$, for $x\in\tau_{t}$, as
$\bigcup_{u<_{\tau}t}\{I(y):y\in^{M}_{u,t}x\}\cup\{\{\tau^{*}_{u}:u<_{\tau}t\}\}.$
@@
Notice that if $\alpha>_{\tau}\beta$ are supertypes, and $x\in\tau^{*}_{\alpha}$, $x\cap\tau^{*}_{\beta}$ is the extension of $x$ over supertype $\beta$. This will be inherited by a scheme of types $\tau_{\gamma}$ with each $\tau_{\gamma}\subseteq\tau^{*}_{\gamma}$ if an additional condition holds: for $\alpha>_{\tau}\beta$, we will have for $x\in\tau_{\alpha}$ that $x\cap\tau^{*}_{\beta}$ is the extension of $x$ over supertype $\beta$ as already noted: for it to be the extension over type $\beta$ we need the general condition $x\cap\tau^{*}_{\delta}\subseteq\tau_{\delta}$ for all $\gamma>\delta$ and $x\in\tau_{\gamma}$.
@@ 3.2 Preliminaries of the construction
On each litter $L=L_{\nu,\beta,\gamma}$ define a well-ordering $\leq_{L}$: $(\nu,\beta,\gamma,\alpha)\leq_{L}(\nu,\beta,\gamma,\alpha^{\prime})$ iff $\alpha\leq\alpha^{\prime}$. The strict well-ordering $<_{L}$ is defined in the obvious way. This well-ordering is used in only one place in the paper (definition 4.2), and its use could easily be avoided, but we find its concreteness appealing.
@@
In our model, any near-litter will be the extension of $\kappa$-amorphous sets in each type above the base type, and we will use these amorphous sets to limit how much the model knows about the relationship between the different extensions of a model element.
@@
A near-litter is determined as the symmetric difference of a litter $L$ (there are $\mu$ litters) and a small subset of $\tau_{-1}$. So it is sufficient to show that a set of size $\mu$ has no more than $\mu$ small subsets. As $\mu$ has cofinality at least $\kappa$, each small subset of $\mu$ has a strict upper bound $\nu<\mu$, and so is an element of $\bigcup_{\nu<\mu}\mathcal{P}(\nu)$. But as $\mu$ is a strong limit cardinal, each $\mathcal{P}(\nu)$ has size less than $\mu$, so there are only $\mu$ bounded subsets of $\mu$; in particular, there are only $\mu$ small subsets of $\mu$. ∎